OpenStudy (anonymous):

Describe a real life scenario where inequalities are used and state the inequality

OpenStudy (anonymous):

you have $20 to buy some cans of soda that cost $.85 each. if we put \(x\) as the number of cans of soda you can buy, then the most you can buy is \[.85x\leq 20\]

OpenStudy (anonymous):

you have $1500 to put as a downpayment on a car, and you have to put at least 20% down. put \(x\) as the price of the car. then he car can cost at most \[1500\geq .20\times x\]